# Time-Transformer: Integrating Local and Global Features for Better Time Series Generation

### 18 Dec 2023 v1

@@ Abstract
Generating time series data is a promising approach to address data deficiency problems. However, it is also challenging due to the complex temporal properties of time series data, including local correlations as well as global dependencies. Most existing generative models have failed to effectively learn both the local and global properties of time series data. To address this open problem, we propose a novel time series generative model named ‘Time-Transformer AAE’, which consists of an adversarial autoencoder (AAE) and a newly designed architecture named ‘Time-Transformer’ within the decoder. The Time-Transformer first simultaneously learns local and global features in a layer-wise parallel design, combining the abilities of Temporal Convolutional Networks and Transformer in extracting local features and global dependencies respectively. Second, a bidirectional cross attention is proposed to provide complementary guidance across the two branches and achieve proper fusion between local and global features. Experimental results demonstrate that our model can outperform existing state-of-the-art models in 5 out of 6 datasets, specifically on those with data containing both global and local properties. Furthermore, we highlight our model’s advantage on handling this kind of data via an artificial dataset. Finally, we show our model’s ability to address a real-world problem: data augmentation to support learning with small datasets and imbalanced datasets.
@@ 1 Introduction
  • Inserting the proposed module into the decoder of theAdversarial Autoencoder, we present a novel time series generative model: Time-Transformer AAE, which efficiently and effectively generates high quality synthetic time series data containing both global and local features.
  • We show empirically that the proposed Time-Transformer AAE can generate better synthetic time series data, with respect to different benchmarks, when compared to SOTA methods.
  • Through comprehensive experiments, we demonstrate the efficiency and effectiveness of our model in handling time series data contains both global and local features, and subsequently, addressing data deficiency issues.
@@ 2 Related Works
2.1 Time Series Generation
@@
Deep generative models (DGMs) have gained increasing attention since their introduction. The variational autoencoder (VAE) uses the Bayesian method to learn latent representations and turn the classic autoencoder into a generative model [28]. Generative adversarial networks (GANs) introduce an adversarial approach to shape the output distribution in order to generate realistic fake data [18]. The adversarial autoencoder (AAE) combines the previous two models together [38]. It uses the adversarial training procedure to perform variational inference in the VAE. Numerous models have been designed based on these basic generative frameworks and shown superior performance in image and text processing [41, 42, 60, 27, 3, 21, 1].
@@
Successes in graphs and text have led to the application of DGMs’ in time series domains. Most of them use the GAN framework with modifications to incorporate temporal properties. The first of these, called C-RNN-GAN [39], directly uses the GAN structure with LSTM to generate music data. Recurrent Conditional GAN (RCGAN) [15] uses basic RNN as backbone, and auxiliary label information as the condition to generate medical time series. Since its introduction, a number of works have utilized similar designs to generate time series data in various fields including finance, medicine and the internet [20, 30, 51, 48, 36]. TimeGAN [57] introduces additional embedding function and supervised loss to generate universal time series. Real-world Time Series GAN (RTSGAN) [43] combines WGAN [3] and autoencoder, and achieves good performance on real-world data generation. PSA-GAN [25] utilizes the progressively growing architecture [27] to generate long time series. To generate both regular and irregular sampled time series, GT-GAN [26] combines Neural Differential Equations and DGMs together for the first time. COSCI-GAN [47] focuses on multivariate time series, and uses a central discriminator to learn better inter-channel dynamics. On the other hand, VAE has also been used for this task. [16] designed a recurrent VAE to synthesize time series data. TimeVAE [13] implements an intepretable temporal structure together with VAE, and achieves SOTA results on universal time series generation. CR-VAE [33] incorporates the causal mechanism into time series generation by learning a Granger causal graph. Recently, the diffusion model has also been utilized together with TCNs to generate audio data [29]. Instead of traditional DGMs, TimeGCI uses the contrastive learning framework to generate time series and achieves good results [24]. These works learn the overall distribution of the training set and generate synthetic time series within the same distribution.
@@
On the other hand, another line of works focus on step-wise generation: They build generative models to infer each time step on the previous ones. Some early works modify the basic RNN using variational inference and get good results with regard to log-likelihood [9, 17]. Some later works combine this idea with TCN [32, 2]. In recent years, diffusion models are used to achieve SOTA results in this area [45, 34]. Though different to our task, these works provide an instructive direction for applying generative models on time series forecasting task.
2.2 Temporal Convolutional Networks and Transformer
@@ 3 Problem Formulation
Generally, multivariate time series generation via GAN-like models involves training a model to approximate a function $\mathcal{F}:\mathcal{P}\rightarrow\mathcal{P}_{d}$ which maps an arbitrary prior distribution $\mathcal{P}$ to the real data distribution $\mathcal{P}_{d}$, so that the model can produce realistic synthetic data $X^{\prime}\sim\mathcal{P}_{d}$ based on any samples $s$ drawn from $\mathcal{P}$. For time series data $X\in\mathbb{R}^{T\times c}$, where $T$ is maximum observable time and $c$ is number of channels. We define the local patterns and the global dependencies as:
@@
where $1<k<t$ and $p$ is the density. They are derived from previous researches [57, 55] with adaptations to align with our requirements. Thus, $X$ contains both $\mathcal{L},\mathcal{G}$ and function $\mathcal{F}$ must learn them simultaneously:
@@ 4.1 Overview
Here, we use Gaussian as the prior distribution $\mathcal{P}=\mathcal{N}(0,1)$ and insert the Time-Transformer after a De-Convolutional block to approximate the $\mathcal{F}$ in (3.3) (This design is to achieve an optimal equilibrium between efficiency and effectiveness, we provide details in D.6). The De-Convolutional block will first reconstruct a prototype of the time series, and then the Time-Transformer aims at learning the detailed local/global features through feedback from reconstruction loss during training. Finally, the model will be able to synthesize realistic data in generation step.
@@ 4.2 Time-Transformer
Layer-wise Parallelization: The Mobile-Former [8] combines MobileNets (a light-weight CNN) and Transformer in parallel, and achieve better performance than sequentially combined models on image classification and object detection tasks. Inspired by this, we also combine TCN and Transformer in a parallel manner to learn $\mathcal{L}$ and $\mathcal{G}$ in (3.3) simultaneously. However, instead of simply combining the entire TCN and Transformer, we only use one layer from the TCN and one block from the Transformer in each Time-Transformer block.
@@
Bidirectional Cross Attention: The bidirectional cross attention block contains two cross attention mechanisms, aims to build interactions between two parallel branches, and thus fuse local and global features, it functions as the $\mathcal{I}$ in (3.3). As illustrated on the right hand side of Figure 2, the output of the TCN layer $\widetilde{\mathcal{L}_{i}}\in\mathbb{R}^{t\times c}$ and the output of the Transformer block $\widetilde{\mathcal{G}_{i}}\in\mathbb{R}^{t\times c}$ interact in a mutual manner within the cross attention block, where it bidirectionally fuses the local feature $\widetilde{\mathcal{L}_{i}}$ and the global feature $\widetilde{\mathcal{G}_{i}}$. Specifically, the TCN features are updated through a residual connection with the attention matrix to obtain $\mathcal{L}_{i+1}$:
@@
where $W_{dq}$, $W_{dk}$, and $W_{dv}$ are learnable parameters of three linear projection layers. $\mathcal{A}_{\widetilde{\mathcal{L}_{i}}\rightarrow\widetilde{\mathcal{G}_{i}}}$ is the affinity matrix from TCN to Trans.
@@ 5.1 Datasets
Sine_Cpx: Another 5000 samples of more complex 5-dimensional sinusoidal sequences. They are simulated using sum of standard sinusoidal waves $x_{t}^{(d)}=\alpha_{1}sin(2\pi f_{1}t+\varphi_{1})+\alpha_{2}sin(2\pi f_{2}t+\varphi_{2})+\alpha_{3}sin(2\pi f_{3}t+\varphi_{3})$ where $\alpha_{i}\sim\mathcal{U}[1,3],f_{i}\sim\mathcal{U}[0.1,0.15],\varphi_{i}\sim\mathcal{U}[0,2\pi],i=1,2,3$ and $\alpha_{i}\neq\alpha_{j},f_{i}\neq f_{j},\varphi_{i}\neq\varphi_{j}$, when $i\neq j$. Thus, the sequences demonstrate global sine waves and local patterns in each dimension.
@@
ECochG: A medical dataset of historical patient data from cochlear implant surgeries [52]. Each instance contains a univariate time series that represents a patient’s inner ear response during the surgery. In general, each time series is a stochastic trend globally, with some local drops. The details of this dataset can be found in A.
5.2 Baseline Models & Benchmarks
We select previous models: TimeVAE, RTSGAN, and TimeGAN [13, 43, 57] as baselines based on their relevance, accessibility of the code, executability of the code, and performance of the model [^1] Models aligned with our task but failed any of these criteria are excluded. E.g., PSA-GAN’s released codes are not executable and RSGAN performs poorly comparing to other baselines. We use their original settings as they have already been optimised by their authors. Details of our model’s settings (hyper-parameters and data pre-processing) can be found in C. To evaluate these generative models, we use a range of metrics and consider several desiderata for the generated time series: They should (1) be distributed close to the original data and different from each other, (2) be indistinguishable from the original data, and (3) preserve the original temporal properties in order to be used for downstream tasks:
@@ 5.2 Baseline Models & Benchmarks
Predictive Score reflects how well the generated data inherit the predictive properties of the original data. We train a post-hoc sequential predictor (a 2-layer LSTM) to do one-more-step prediction, and calculate the mean absolute error (MAE). Here, we use the ‘Train on Synthetic, Test on Real’ (TSTR) technique introduced by [15]. The predictor should be able to learn from the generation and predict the real data. The Predictive Score provides a quantitative assessment of (3).
@@ 5.3 Experimental Results
5.3.1 Visualisation & FID
Figure 3 shows visualization results of three local-global datasets (t-SNE plots for preliminary datasets can be found in D.1). Each row contains plots of one dataset and each column represents a model. The blue and red dots represent original data points and generations respectively. It is obvious that our proposed Time-Transformer AAE (denoted as Time-Trans. hereafter) consistently produces synthetic data closely distributed to the original data. TimeVAE and RTSGAN also have good distributions for most datasets. Table 1 shows the FID scores of each model on each dataset. Our proposed Time-Trans. obtains the best scores for all the datasets except ‘Stock’ where RTSGAN is slightly better. For those datasets where the models show similar visualisation results (e.g. Time-Trans. and TimeVAE on ‘Sine_Cpx’), we can use the FID scores to quantitatively evaluate them.
@@ 5.3.1 Visualisation & FID
5.3.2 Discriminative Score & Predictive Score
As per the Discriminative Scores shown in Table 2, our Time-Trans. performs well for most datasets, and particularly excels in performance for the local-global datasets where the generations from our model show better indistinguishability than the other methods.
@@ 5.3.2 Discriminative Score & Predictive Score
We report ten times MAE as Predictive Score to avoid long decimal numbers while highlighting significant differences. An additional column ‘Oracle’ shows results from the ‘Train on Real, Test on Real’, which represent the theoretically best performances. Table 3 shows a similar trend: our model performs well in most datasets, particularly, in local-global datasets where our model’s scores are close to the oracle. This indicates that our model has learnt most of the predictive properties of these data. Note that scores in our work differ to those published in related works because we calculate them in a slightly different way. We provide a detailed justification in D.9. Apart from these six datasets, we also evaluated our model on longer time series data. The results show that our model consistently out-performs baselines with these longer time series (details can be found in D.2).
5.3.3 Local & Global Features Learning
@@
The green and the orange series represent local and global features respectively. They are generated with high (50Hz) and low (5Hz) frequencies with respect to the Fast Fourier Transform (FFT). By adding them together we get the blue series which, contains both global and local features (named ‘Mixture’). The formulae to generate local, global and mixture time series are alongside, where $\alpha\sim\mathcal{U}[1,3]$.
@@
We vary the portion of each part in Mixture to get two other series that contain 75% local and 75% global respectively. We train the generative models on these datasets and compare performance. Figure 5 shows the above scores (mean score with standard deviation) of the three best models (RTSGAN, TimeVAE, and Time-Trans.).
Figure 5: Results of Time-Trans.(green), RTSGAN(blue), and TimeVAE(orange). Lower are better.
When the data only contains one feature type, all models can generate comparatively good time series, but their performances decrease with mixed features. Among these models, ours shows the most consistent performance in all cases and has the best performance in mixture datasets with respect to all metrics, indicating its effectiveness in learning both local and global features. We provide detailed results in D.3.
5.3.4 Model Application: Data Augmentation
We evaluate our model on imbalanced classification using the ECochG dataset and three real-world datasets from the UCR archive [11]: Wafer, Herring, and SwedishLeaf (details in B). Previous studies argued oversampling approaches (via replication or simple modifications like jittering) can cause overfitting or create out-of-distribution synthesis [56, 54]. Our model is expected to address these issues. Table 4 lists the classification results on the ECochG dataset using an off-the-shelf Multi-layer Perceptron (MLP) with (1) no augmentations, and augmenting the minority class via (2) repeating (3) jittering, (4) RTSGAN, (5) TimeVAE, and (6) our model (augmentations are only in training sets).

### 28 Jan 2024 v3

@@ Abstract
Generating time series data is a promising approach to address data deficiency problems. However, it is also challenging due to the complex temporal properties of time series data, including local correlations as well as global dependencies. Most existing generative models have failed to effectively learn both the local and global properties of time series data. To address this open problem, we propose a novel time series generative model named ‘Time-Transformer AAE’, which consists of an adversarial autoencoder (AAE) and a newly designed architecture named ‘Time-Transformer’ within the decoder. The Time-Transformer first simultaneously learns local and global features in a layer-wise parallel design, combining the abilities of Temporal Convolutional Networks and Transformer in extracting local features and global dependencies respectively. Second, a bidirectional cross attention is proposed to provide complementary guidance across the two branches and achieve proper fusion between local and global features. Experimental results demonstrate that our model can outperform existing state-of-the-art models in 5 out of 6 datasets, specifically on those with data containing both global and local properties. Furthermore, we highlight our model’s ability to handle this kind of data via an artificial dataset. Finally, we show how our model performs when applied to a real-world problem: data augmentation to support learning with small datasets and imbalanced datasets.
@@ 1 Introduction
  • By inserting the proposed module into the decoder of the Adversarial Autoencoder, we present a novel time series generative model, Time-Transformer AAE, which efficiently and effectively generates high quality synthetic time series data containing both global and local features.
  • We show empirically that the proposed Time-Transformer AAE can generate better synthetic time series data with respect to different benchmarks, when compared to SOTA methods.
  • Through comprehensive experiments, we demonstrate the efficiency and effectiveness of our model in handling time series data containing both global and local features, and subsequently, addressing data deficiency issues.
@@ 2 Related Works
2.1 Time Series Generation:
@@
Deep generative models (DGMs) have gained increasing attention since their introduction. The variational autoencoder (VAE) uses the Bayesian method to learn latent representations and turn the classic autoencoder into a generative model [28]. Generative adversarial networks (GANs) introduce an adversarial approach to shape the output distribution in order to generate realistic fake data [18]. The adversarial autoencoder (AAE) combines the previous two models together [39]. It uses the adversarial training procedure to perform variational inference in the VAE. Numerous models have been designed based on these basic generative frameworks that have shown superior performance in image and text processing [42, 43, 61, 27, 3, 21, 1].
@@
Successes in processing graphs and text have led to the application of DGMs’ in time series problems. Most of them use the GAN framework with modifications to incorporate temporal properties. The first of these, called C-RNN-GAN [40], directly uses the GAN structure with LSTM to generate music data. Recurrent Conditional GAN (RCGAN) [15] uses a basic RNN as its backbone, and auxiliary label information as the condition to generate medical time series. Since its introduction, a number of works have utilized similar designs to generate time series data in various fields including finance, medicine and the internet [20, 30, 52, 49, 36]. TimeGAN [58] introduces additional embedding function and supervised loss to generate universal time series. Real-world Time Series GAN (RTSGAN) [44] combines WGAN [3] and autoencoder, and achieves good performance on real-world data generation. PSA-GAN [25] utilizes the progressively growing architecture [27] to generate long time series. To generate both regular and irregular sampled time series, GT-GAN [26] combines Neural Differential Equations and DGMs together for the first time. COSCI-GAN [48] focuses on multivariate time series, and uses a central discriminator to learn better inter-channel dynamics. On the other hand, VAE has also been used for this task. [16] designed a recurrent VAE to synthesize time series data. TimeVAE [13] implements an intepretable temporal structure together with VAE, and achieves SOTA results on universal time series generation. CR-VAE [33] incorporates the causal mechanism into time series generation by learning a Granger causal graph. Recently, the diffusion model has also been utilized together with TCNs to generate audio data [29]. In comparison to traditional DGMs, TimeGCI uses the contrastive learning framework to generate time series and achieves good results [24]. These works learn the overall distribution of the training set and generate synthetic time series within the same distribution.
@@
On the other hand, another line of works focus on step-wise generation: They build generative models to infer each time step on the previous ones. Some early works modify the basic RNN using variational inference and achieve good results with regard to log-likelihood [9, 17]. Some later works combine this idea with TCN [32, 2]. In recent years, diffusion models are used to achieve SOTA results in this area [46, 34]. Though different to our task, these works provide an instructive direction for applying generative models on the time series forecasting task.
2.2 Temporal Convolutional Networks and Transformer:
@@ 3 Problem Formulation
Generally, multivariate time series generation via GAN-like models involves training a model to approximate a function $\mathcal{F}:\mathcal{P}\rightarrow\mathcal{P}_{d}$ which maps an arbitrary prior distribution $\mathcal{P}$ to the real data distribution $\mathcal{P}_{d}$, so that the model can produce realistic synthetic data $X^{\prime}\sim\mathcal{P}_{d}$ based on any samples $s$ drawn from $\mathcal{P}$. For time series data $X\in\mathbb{R}^{T\times c}$, where $T$ is maximum observable time and $c$ is number of channels, we define the local patterns and the global dependencies as:
@@
where $1<k<t$ and $p$ is the density. They are derived from previous researches [58, 56] with adaptations to align with our requirements. Thus, $X$ contains both $\mathcal{L}and\mathcal{G}$ and function $\mathcal{F}$ must learn them simultaneously:
@@ 4.1 Overview:
Here, we use Gaussian as the prior distribution $\mathcal{P}=\mathcal{N}(0,1)$ and insert the Time-Transformer after a De-Convolutional block to approximate the $\mathcal{F}$ in (3.3) (This design is to achieve an optimal equilibrium between efficiency and effectiveness. We provide details in [38]). The De-Convolutional block will first reconstruct a prototype of the time series, and then the Time-Transformer aims at learning the detailed local/global features through feedback from reconstruction loss during training. Finally, the model will be able to synthesize realistic data in the generation step.
@@ 4.2 Time-Transformer:
Layer-wise Parallelization: The Mobile-Former [8] combines MobileNet (a light-weight CNN) and Transformer in parallel, and achieve better performance than sequentially combined models on image classification and object detection tasks. Inspired by this, we also combine TCN and Transformer in a parallel manner to learn $\mathcal{L}$ and $\mathcal{G}$ in (3.3) simultaneously. However, instead of simply combining the entire TCN and Transformer, we only use one layer from the TCN and one block from the Transformer in each Time-Transformer block.
@@
Bidirectional Cross Attention: The bidirectional cross attention block contains two cross attention mechanisms, aims to build interactions between two parallel branches, and thus fuses local and global features to function as the $\mathcal{I}$ in (3.3). As illustrated on the right hand side of Figure 2, the output of the TCN layer $\widetilde{\mathcal{L}_{i}}\in\mathbb{R}^{t\times c}$ and the output of the Transformer block $\widetilde{\mathcal{G}_{i}}\in\mathbb{R}^{t\times c}$ interact in a mutual manner within the cross attention block, where it bidirectionally fuses the local feature $\widetilde{\mathcal{L}_{i}}$ and the global feature $\widetilde{\mathcal{G}_{i}}$. Specifically, the TCN features are updated through a residual connection with the attention matrix to obtain $\mathcal{L}_{i+1}$:
@@
where $W_{dq}$, $W_{dk}$, and $W_{dv}$ are learnable parameters of three linear projection layers. $\mathcal{A}_{\widetilde{\mathcal{L}_{i}}\rightarrow\widetilde{\mathcal{G}_{i}}}$ is the affinity matrix from TCN to Trans
@@ 5.1 Datasets:
Sine_Cpx: Another 5000 samples of more complex 5-dimensional sinusoidal sequences. They are simulated using the sum of standard sinusoidal waves $x_{t}^{(d)}=\alpha_{1}sin(2\pi f_{1}t+\varphi_{1})+\alpha_{2}sin(2\pi f_{2}t+\varphi_{2})+\alpha_{3}sin(2\pi f_{3}t+\varphi_{3})$ where $\alpha_{i}\sim\mathcal{U}[1,3],f_{i}\sim\mathcal{U}[0.1,0.15],\varphi_{i}\sim\mathcal{U}[0,2\pi],i=1,2,3$ and $\alpha_{i}\neq\alpha_{j},f_{i}\neq f_{j},\varphi_{i}\neq\varphi_{j}$, when $i\neq j$. Thus, the sequences demonstrate global sine waves and local patterns in each dimension.
@@
ECochG: A medical dataset of historical patient data from cochlear implant surgeries [53]. Each instance contains a univariate time series that represents a patient’s inner ear response during the surgery. In general, each time series is a stochastic trend globally, with some local drops. More details of this dataset can be found in [38].
5.2 Baseline Models & Benchmarks:
We select previous models, TimeVAE, RTSGAN, and TimeGAN [13, 44, 58], as baselines based on their relevance, accessibility of the code, executability of the code, and performance of the model [^1] Models aligned with our task but failed any of these criteria are excluded. E.g., PSA-GAN’s released codes are not executable. We use their original settings as they have already been optimised by their authors. Details of our model [^2] The code is available at https://github.com/Lysarthas/Time-Transformer (hyper-parameters and data pre-processing) can be found in [38]. To evaluate these generative models, we use a range of metrics and consider several desiderata for the generated time series: They should (1) be distributed close to the original data and different from each other, (2) be indistinguishable from the original data, and (3) preserve the original temporal properties in order to be used for downstream tasks.
@@ 5.2 Baseline Models & Benchmarks:
Predictive Score reflects how well the generated data inherits the predictive properties of the original data. We train a post-hoc sequential predictor (a 2-layer LSTM) to do one-more-step prediction, and calculate the mean absolute error (MAE). Here, we use the ‘Train on Synthetic, Test on Real’ (TSTR) technique introduced by [15]. The predictor should be able to learn from the generation and predict the real data. The Predictive Score provides a quantitative assessment of (3).
@@ 5.3 Experimental Results
5.3.1 Visualisation & FID:
Figure 3 shows visualization results of three local-global datasets (t-SNE plots for preliminary datasets can be found in [38]). Each row contains plots of one dataset and each column represents a model. The blue and red dots represent original and generated data points respectively. It is obvious that our proposed Time-Transformer AAE (denoted as Time-Trans hereafter) consistently produces synthetic data closely distributed to the original data. TimeVAE and RTSGAN also have good distributions for most datasets. Table 1 shows the FID scores of each model on each dataset. Our proposed Time-Trans obtains the best scores for all the datasets except ‘Stock’ where RTSGAN is slightly better. For those datasets where the models show similar visualisation results (e.g. Time-Trans and TimeVAE on ‘Sine_Cpx’), we can use the FID scores to quantitatively evaluate them.
@@ 5.3.1 Visualisation & FID:
5.3.2 Discriminative Score & Predictive Score:
As per the Discriminative Scores shown in Table 2,
@@ 5.3.2 Discriminative Score & Predictive Score:
our Time-Trans performs well for most datasets, and particularly excels in performance for the local-global datasets where the generations from our model show better indistinguishability than the other methods.
We report ten times MAE as Predictive Score to avoid long decimal numbers while highlighting significant differences. An additional column ‘Oracle’ shows results from the ‘Train on Real, Test on Real’, which represent the theoretically best performances. Table 3 shows a similar trend: our model performs well in most datasets, particularly, in local-global datasets where our model’s scores are close to Oracle. This indicates that our model has learnt most of the predictive properties of these datasets. Note that scores in our work differ to those published in related works because we calculate them in a slightly different way. We provide a detailed justification in [38]. Apart from these six datasets, we also evaluated our model on longer time series data. The results show that our model consistently out-performs baselines with these longer time series (details can be found in [38]).
5.3.3 Local & Global Features Learning:
@@
The green and orange series represent local and global features respectively. They are generated with high (50Hz) and low (5Hz) frequencies with respect to the Fast Fourier Transform (FFT). By adding them together we get the blue series which, contains both global and local features (named ‘Mixture’). The formulae to generate local, global and mixture time series are alongside, where $\alpha\sim\mathcal{U}[1,3]$.
@@
We vary the portion of each part in Mixture to get two other series that contain 75% local and 75% global respectively. We train the generative models on these datasets and compare performance. Figure 5 shows the above scores (mean score with standard deviation) of the three best models (RTSGAN, TimeVAE, and Time-Trans).
Figure 5: Results of Time-Trans(green), RTSGAN(blue), and TimeVAE(orange). Lower are better.
When the data only contains one feature type, all models can generate comparatively good time series, but their performances decrease with mixed features. Among these models, ours shows the most consistent performance in all cases and has the best performance in mixture datasets with respect to all metrics, indicating its effectiveness in learning both local and global features. We provide detailed results in [38].
5.3.4 Model Application - Data Augmentation:
We evaluate our model on imbalanced classification using the ECochG dataset and three real-world datasets from the UCR archive [11]: Wafer, Herring, and SwedishLeaf (details in [38]). Previous studies argued oversampling approaches (via replication or simple modifications like jittering) can cause overfitting or create out-of-distribution synthesis [57, 55]. Our model is expected to address these issues. Table 4 lists the classification results on the ECochG dataset using an off-the-shelf Multi-layer Perceptron (MLP) with (1) no augmentations, and augmenting the minority class via (2) repeating (3) jittering, (4) RTSGAN, (5) TimeVAE, and (6) our model (augmentations are only in training sets).